July 13, 2016 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, July 13, 2016
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or bjn.vc
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2016/ .

Calibration Tasks

  1. Spring 2016 Recon Checklist
  2. Mcsmear updates - mcsmear 20160323
  3. Data Validation
  4. GlueX Detector Performance Document

Calibrations for Second Production Run

  • Updated error matrices (all)
  • Fix handling of individually-read-out TAGM columns (Alex B.)
  • Tagger hit merging (Nathan/Alex B.)
  • Updates to FCAL non-linearity (Adesh)
  • Recalibrate gains for early runs ~<10600 (Adesh)
  • Add timing information to FCAL clustering/inner ring calibrations (Adesh/Matt S.)
  • BCAL TDC calibrations (Andrei)
  • BCAL energy non-linearity correction (Will M.)
  • FDC wire timing & reconstruction (Lubomir/Simon/Alex A.)
  • TOF ADC energy and timing calibrations (Beni)
  • TOF alignment (Beni/Simon)
  • General timing realignment (Sean)


  1. Announcements
  2. Calibration Plans for Second Production Run
    • Calorimetry
    • Tracking
    • Start Counter
    • Photon Beam
  3. Next Monitoring Run?
  4. Simulations
    • sim1.1 for data/simulation matching studies (Sean/Mark I.)
    • Mcsmear updates
      • Detector updates
  5. AOB


Attending: Sean (NU); Paul M., Simon, Adesh, Mark D., Justin, Alex A., Alex B., Nathan, Elton, Mark D., Lubomir (JLab); Curtis, Mike S., Will M. (CMU); Mahmoud (FIU)

  1. Announcements
    • At yesterday's Analysis Meeting, Elton showed some interesting results which (among other things) indicated that the tracking errors going into the kinematic fit needed improvement. Simon is working on this, and has already made some progress. Also, Mahmoud's Start Counter efficiency measurements may have some biases. Paul had some suggestions for Mahmoud - these and other plans will be discussed at tomorrow's SC meeting.
  2. Calibration Plans for Second Production Run
    • FCAL - Adesh is almost finished with the timing calibrations. He has also been looking at the results from the last reconstruction launch. Updates and lessons learned will be presented at tomorrow's Calorimetry meeting.
    • BCAL - Will has been looking at data/MC comparisons to see why the pi0 resolution differs. He thinks the sampling fluctuations and timing smearing are likely at fault, and is looking into improving these. The next step will be to run the gain calibration procedure on the simulated data to see what calibrations come out.
    • CDC - Mike presented several works-in-progress at the Analysis meeting yesterday. Data/MC matching is in progress.
    • FDC - Progress is being made on several fronts. Lubomir (using GenFit) and Simon (default sim-recon Kalman filter) have been comparing results with two different tracking code. They have tracked down a few sign errors in how the alignment constants were being applied. They studied the drift time/distance relation and found that what is seen in data is different from the Garfield calculation, but Garfield does get the with/without B-field scaling correct (~10%). The Lorentz deflections are a larger effect, and they are under control. Simon has included these new distances into code in a private branch. Alex A. presented results yesterday about his improvements on FDC pseudo-hit reconstruction and updates to the timing calibrations.
    • TOF - Beni: No news is good news.
    • Start Counter - In looking at three runs along the spring, Mahmoud sees a ~20 ps change in time resolutions. He is going to do a more systematic study.
    • TAGM - Alex B. committed code to sim-recon that disables the individually-readout columns from being used in reconstruction, which was giving some double-counting. He is working on clustering hits from photons that hit multiple fibers.
    • TAGH - Nathan is working on adding hit merging, which is giving a few percent effect. ~10% of the merged hits are due to accidentals, which is overall a few-tenths-of-a-percent effect. There are some minor overlaps between the TAGH and TAGM, Nathan will put together a description of what he is seeing.
  3. With the improvements in reconstruction, especially the FDC, it would be good to have another monitoring launch. Simon and Alex A. think that their improvements should be ready by the end of the week. Paul agrees that a launch can be scheduled for the end of the month.
  4. Simulations
    • Mark I. has updated the plugin list and populated the jobs table with events from different runs, in proportion to the number of triggers in each run. He is using RCDB in his scripts to decide which collimator to use - the affects the photon spectrum used in bggen and hdgeant. He is running some final checks before turning everything on. Sean mentioned that based on previous discussions, it would be good to start
    • Sean is doing some final tests of his changes to mcsmear. The changes include realistic time resolutions for TOF and SC, machinery for per-channel efficiency corrections for data/MC matching, and machinery for extracting run configuration info from RCDB. Currently, people are looking at the BCAL, FCAL, and FDC smearing parameterizations. More attention is needed for the CDC, TOF, and SC.